# On the origins of counterrotating stellar disks in TNG50. I

### v1

@@ Abstract
Context. Understanding galaxy evolution is key to explaining the structures we observe today in the present-day Universe. Counterrotating stellar disks, i.e. co-spatial stellar disks rotating with opposite angular momentum, have been proposed as signatures of past accretion events. Therefore, they constitute potentially valuable tracers of galactic assembly.
Aims. In this work, we aim to investigate the properties, formation channels, and significance of counterrotating stellar disks in a sample of Milky Way mass galaxies using the IllustrisTNG cosmological simulations.
Methods. We select an initial sample of 260 central late-type galaxies (i.e. $M_{\rm tot}\approx 10^{12}$, $D/T>0.5$, $N_{\rm star}>10^{5}$). For each galaxy, we measure the circularity of its stellar particles and we define the counterrotating disk by considering all stellar particles with circularity $\epsilon<-0.7$, which are located within the spatial extension of the main disk. We then characterize the mass fraction, spatial extent, and star formation history of the counterrotating disks.
Results. Out of the 260 central late-type galaxies, we find that 26 host significant counterrotating disks (i.e. contributing at least 1% of the total stellar mass of the disk). This means that counterrotating stellar disks (CRDs) are rare, consistent with the results from recent observations. We also find that the most of the CRDs are compact (i.e. 88%), in-situ dominated (i.e.73%), and exhibit bursty SFHs whose peaks often coincide with external perturbations. This means that external perturbations are able to catalyze retrograde star formation, even when a majority of the CRD’s star population is in-situ. Finally, we find that a variety of formation pathways can lead to CRDs, including interaction-induced in-situ bursts and smooth accretion of misaligned gas. This fuels star formation in the CRD component
Conclusions. Overall, our results suggest that CRDs are rare but diverse in origin. In most cases, their formation is linked to the accretion of retrograde gas, either through mergers or environmental inflow, suggesting that these structures are sensitive tracers of the galaxy’s past accretion history.
@@ 1 Introduction
Present-day galaxies exhibit a wide variety of structural and kinematic properties (e.g., Monachesi et al. 2016; Rowlands et al. 2018), shaped by both internal and external processes such as minor mergers, ram-pressure stripping, and galaxy harassment (Boselli & Gavazzi, 2006). Each of these mechanisms leaves long lasting features that can be used to characterize their evolution. A particularly revealing class of features are multi-spin components (Rubin, 1994), that is, kinematic stellar structures whose angular momentum is misaligned with respect to the main stellar disk. These components highlight the complexity of galactic dynamics, as present-day kinematics preserve the imprints of past accretion and interaction events (Monachesi et al. 2016; Rowlands et al. 2018; Monachesi et al. 2019; Dolfi et al. 2023). Among them, co-spatial counterrotating stellar disks (Braun et al., 1994) are of particular interest. These consist of two co-existing stellar components that are strongly misaligned in angular momentum but confined to the same plane. Their formation pathways and demographics provide unique constraints on the role of external accretion in shaping galaxy evolution.
@@
The presence of counterrotating stellar disks (hereafter CRDs) is noteworthy because their high velocity amplitudes make it possible to distinguish their stellar populations from those of the main disk, thereby providing insights into the formation and evolution of their host galaxies. It has been argued that CRDs originate from the retrograde acquisition of gas, which settles on misaligned orbits with respect to the pre-existing stellar disk and subsequently forms stars (Pizzella et al. 2004; Roškar et al. 2010; van de Voort et al. 2015; Starkenburg et al. 2019; Khoperskov et al. 2021, Bao et al. 2022; Bao et al. 2024; Gasymov et al. 2025). CRDs are therefore direct evidence of past accretion events and constitute valuable tracers of galactic assembly. Several mechanisms have been proposed to explain the origin of the retrograde gas, including accretion from the cosmic web (Algorry et al. 2014), gas-rich mergers with dwarf galaxies (Thakar & Ryden 1996; Puerari & Pfenniger 2001; Algorry et al. 2014), tidal gas exchange between neighboring galaxies (Khim et al. 2021; Sil’chenko et al. 2023), and major mergers that preserve the stellar disk (Puerari & Pfenniger 2001; Crocker et al. 2009). Among these, the latter appears to be the least common in observed systems.
@@
Observationally, CRDs appear to be rare, with only 1% of late-type galaxies presenting a CRD (Bao et al. 2022, Bevacqua et al. 2022, Gasymov et al. 2025). At the same time, CRDs are also challenging to observe due to their low mass contribution (Rubino et al. 2021). Nevertheless, modern Integral Field Spectroscopy (IFS) instruments can be used to observe CRDs, which are typically identified from the presence of bimodal features in the velocity dispersion maps (Bao et al. 2022; Bao et al. 2024; Gasymov et al. 2025). IFS instruments are also essential to disentangle the two stellar disk components and study in detail the kinematics and stellar population properties of CRDs, allowing us to constrain their formation scenarios. Complementary techniques, such as SED fitting, are also used to extract physical and chemical properties of the CRDs. Some notable examples of CRDs include: NGC 4550 (Rubin et al. 1992; Rix et al. 1992; Coccato et al. 2013), NGC 7217 (Merrifield & Kuijken, 1994), NGC 3593 (Bertola et al. 1996; Corsini et al. 1998; García-Burillo et al. 2000; Coccato et al. 2013), NGC 4138 (Jore et al., 1996) and NGC 5719 (Vergani et al. 2007; Coccato et al. 2011). These CRDs show a broad range of properties. They are generally observed to have younger stellar populations that co-rotate with the gas disk (Pizzella et al. 2004; Pizzella et al. 2014; Bao et al. 2022; Gasymov et al. 2025), steeper metallicity gradients and broader metallicity distributions than the main stellar disk (Coccato et al. 2013; Bevacqua et al. 2022; Gasymov et al. 2025). As previously mentioned, the presence of CRDs is typically linked to past accretion events that induce SF within the galaxy. Previous simulations (e.g. TNG100, EAGLE etc) have begun to explore the formation and evolution of CRD components (Lu et al. 2021; Santucci et al. 2024; Peirani et al. 2025). However, it is still unclear which are the dominant CRD formation mechanisms and under what conditions CRDs arise, particularly in late-type galaxies, which remains under-explored compared to the early-type galaxy counterparts, that are more common (Kuijken et al., 1996).
@@
In this work we focus on the search, characterization and analysis of counter-rotating stellar disks (CRDs, hereafter) in central Milky Way (MW) type galaxies ($M_{\rm tot}\approx 10^{12}M_{\odot}$) in TNG50 from the IllustrisTNG suite of cosmological hydrodynamical simulations. We aim to determine how common CRDs are in MW-type galaxies, as well as what are the dominant CRD formation mechanisms. In particular, we aim to confirm past formation channels proposed by observations of CRDs. The paper is organized as follows. In Section 2 we give an overview of the TNG50 simulation that is used, as well as define our sample of galaxies, and our definition of counterrotation. In Section 3, an initial characterization of counterrotating components in the galaxies is shown, afterwards, the selection of significant CRDs considered is defined. In Section 4, the structural properties and assembly of the significant CRDs is shown, ending with a classification of the CRDs. We analyze the possible formation channels for the CRDs in each classification in Section 5. We discuss our results in Section 6. Finally, we make a summary of our obtained results and conclude in Section 7.
@@ 2 Methodology
In this section we discuss the simulations employed in this paper, as well as the galaxy sample selection criteria and the definition of counterrotating stellar disks.
@@ 2.1 The IllustrisTNG: TNG50
Table 1: Parameters of the sample of galaxies with a significant CRD component, as well as the six boundary cases. Column descriptions from the left to the right are: (1) internal name, (2) z=0 SubfindID from the TNG50 simulations, (3) total gravitationally bounded mass as defined by the Subfind algorithm, (4) stellar mass of the counterrotating disk, (5) total stellar disk mass, (6) half-mass radius of the stellar disk, (7) optical radius, (8) half-mass radius of the counterrotating stellar disk, (9) radius enclosing 95% of the counterrotating stellar disk mass, (10) stellar concentration of the counterrotating disk as defined in Sec. 4, (11) fraction of ex-situ stellar particles in the CRD.
@@
[…]
The TNG50 data provides the positions, velocities, mass, metallicity and metal abundances of the stellar particles as part of the complete snapshot data. Each stellar particle represents a single stellar population, with a given age and metallicity that is hereditary from the progenitor gas cell. Information about the galaxy and time where the stellar particle is born is provided in Rodriguez-Gomez et al. (2015) and is described in more detail in Section 2.3. The TNG50 simulation also gives information about the galaxies extracted from their member particles as a whole ”classified according to a FoF algorithm, and subsequently Subfind”, such as their total mass, total stellar mass, and virial radius. Information about the disk mass fraction according to the circularity of the particles is provided as an additional catalog by Genel et al. (2015).
@@ 2.2 Sample selection
In this study we decided to focus on a sample of MW-type galaxies. In a follow up study we will expand this sample to include a wider range of late-type galaxies. In particular, we select galaxies that fulfill the following criteria:
@@
  • The galaxy must be the central galaxy in its group, meaning that it should be the most submassive halo according to the Friends-of-Friends(FOF) algorithm.
@@
In particular, the D/T ratios are obtained from the publicly available database of IllustrisTNG as part of the additional catalogs111www.tng-project.org/data/docs/specifications/#sec5c, and are computed by considering as disk stellar particles the ones with circularity parameter (see Sec 2.3) larger than 0.7.
@@
After applying the described selection criteria, we obtain a total of 260 late-type disk galaxies. As discussed, selected galaxies have a number of stellar particles of $\leq 5\times 10^{5}$ stellar particles within them, thus ensuring enough particles to reasonably resolve the structure of the stellar disk and search for potential structures with small mass contributions, such as counterrotating stellar components.
@@
Figure 1: Distribution of the counterrotating disk mass fraction for the full sample of 260 late-type disk galaxies. The vertical dashed orange line indicates a 1% mass fraction from the CRD. We see that only 10% of the galaxies (when including the boundary cases) have a CRD mass fraction greater than 1%.
2.3 Counterrotating disk definition
Figure 2: The face-on and edge-on projections of the stellar mass surface density for a sub-sample of 8 galaxies with the most significant counterrotating stellar disks. The galaxies are sorted according to their total mass, from the most to the least massive. Here, we show all the stellar particles that belong to the host galaxy. The yellow solid and dot-dashed circle represent the radii, $R_{95}^{\rm CRD}$ and $R_{50}^{\rm CRD}$, of the counterrotating disk, respectively, while the white dashed circle represents the optical radius, $R_{\rm opt}$, of the galaxy.
Figure 3: Same as Fig. 2, but we show here only the stellar particles belonging to the CRDs.
@@ 2.3 Counterrotating disk definition
The goal of this work is to identify, quantify and analyze counterrotating stellar disks components (CRDs) in fully cosmological simulations of galaxy formation. Different methods for isolating particles associated to different galaxy components, such as counterrotating disks have been proposed. For example, Tissera et al. (2012) applies a criteria based on circularity and binding energy to subdivide a late type galaxy on its different components. In this study, we follow an approach more closely related to the definitions used in Gómez et al. (2017) and Monachesi et al. (2019). We first rotate the galaxy so that the plane of the disk is co-planar with the xy-plane, i.e the angular momentum of the galactic disk is parallel to the Z-axis direction. Afterwards, we identify particles that are counter-rotating in the galaxy using the circularity parameter, thus performing a kinematic decomposition. Following Abadi et al. (2003) this parameter is defined as:
@@
where $J_{z}$ is the angular momentum in the z-axis and the term $\mathrm{max}(J_{z}(E))$” indicates the maximum angular momentum in the z-direction allowed for that orbital energy; i.e. that of a circular orbit.
In order to define the extension of the stellar disks we use the optical radius, $R_{\rm opt}$. As in Varela-Lavin et al. 2023 and Dolfi et al. 2023, $R_{\rm opt}$ is defined as the radius at which the surface-brightness (SB) profile on the V-band of the galaxy, orientated on a face-on projection, reaches $\mu_{V}=25\ \mathrm{mag\ arcsec}^{-2}$. To estimate the optical radius, we measure the surface-brightness of the galaxy in cylindrical shells of width 0.5kpc and height 10kpc above and below the disk plane. The values of the $R_{\rm opt}$ obtained for each galaxy are listed on Table 1.
Having the circularity parameter computed for every stellar particle within $R_{\rm opt}$. We define the counterrotating disk components by selecting stellar particles that fulfill the following criteria:
@@
  • A distance from the galactic plane of under 5kpc ($|z|<5\ \rm kpc$).
  • A galactocentric distance under $R_{\rm opt}$.
These selection criteria allow us to select star particles in nearly circular orbits rotating in the opposite direction of the overall disk.
In this work we are also interested on characterizing the formation mechanisms of CRDs. Due to this, it is important to identify the origins of the different stellar particles associated to these components. In particular, we subdivide the counterrotating stellar particles as in-situ or ex-situ based on where the particle were born. As discussed in Rodriguez-Gomez et al. 2016, a particle is considered in-situ if the galaxy in which it formed is part of the main branch of the merger tree of the galaxy where the particle is at the present-day. In case the particle was born outside this main branch then the particle is tagged as ex-situ. Notice that particles that are formed within the potential well of the main progenitor from recently stripped gas from satellite galaxies are labeled as in-situ.
3 Identification of counterrotating disks

### v2

# Origins of coplanar counterrotating stellar disk components in late-type galaxies. I

@@ Abstract
Context. Understanding galaxy evolution is key to explaining the structures we observe in the present-day Universe. Counterrotating stellar disks, i.e., co-spatial stellar disks rotating with opposite angular momentum, have been proposed as signatures of past accretion events. Therefore, they constitute potential tracers of galactic assembly.
Aims. In this work, we aim to investigate the properties, formation channels, and significance of coplanar counterrotating disk components (CRDs) in a sample of Milky Way mass galaxies using the IllustrisTNG cosmological simulations.
Methods. We selected an initial sample of 260 central late-type galaxies (i.e. $M_{\rm tot}\approx 10^{12}$, $D/T>0.5$, $N_{\rm star}>10^{5}$). For each galaxy, we measured the circularity of its stellar particles, and we defined a CRD by considering all particles with circularity $\epsilon<-0.7$, which are located within the spatial extension of the main disk. We then characterized the mass fraction, spatial extent, and star formation history of the CRDs.
Results. Out of the 260 late-type galaxies, we find 26 host significant CRDs (i.e., contributing at least 1% of the mass of the stellar disk). This means that CRDs are rare, and this outcome is consistent with the results from recent observations. We also find that most of the CRDs are compact (i.e., 88%) and in situ dominated (i.e., 73%), and they exhibit bursty star formation histories whose peaks often coincide with external perturbations. This means that external perturbations are able to catalyze retrograde star formation, even when a majority of the CRD’s star population is in situ. Finally, we find that a variety of formation pathways can lead to CRDs, including interaction-induced in situ bursts and smooth accretion of misaligned gas.
Conclusions. Overall, our results suggest that CRDs are rare but diverse in origin. In most cases, their formation is linked to the accretion of retrograde gas, either through mergers or environmental inflow, suggesting that these are sensitive tracers of the galaxy’s past accretion history.
@@ 1 Introduction
Present-day galaxies exhibit a wide variety of structural and kinematic properties (e.g., Monachesi et al. 2016; Rowlands et al. 2018) shaped by both internal and external processes such as minor mergers, ram-pressure stripping, and galaxy harassment (Boselli and Gavazzi, 2006). Each of these mechanisms leaves long-lasting features that can be used to characterize their evolution. A particularly revealing class of features are multi-spin components (Rubin, 1994), that is, kinematic stellar structures whose angular momentum is misaligned with respect to the main stellar disk. These components highlight the complexity of galactic dynamics, as present-day kinematics preserve the imprints of past accretion and interaction events (Monachesi et al. 2016; Rowlands et al. 2018; Monachesi et al. 2019; Dolfi et al. 2023). Among these misaligned components, co-spatial counterrotating stellar disks (Braun et al., 1994) are of particular interest. They consist of two coexisting stellar components that are strongly misaligned in angular momentum but confined to the same plane. Their formation pathways and demographics provide unique constraints on the role of external accretion in shaping galaxy evolution.
@@
The presence of coplanar counterrotating stellar disk components (CRDs) is noteworthy because, if their mass contribution is significant enough, their high velocity amplitudes make it possible to distinguish their stellar populations from those of the main disk, thereby providing insights into the formation and evolution of their host galaxies. It has been argued that counterrotating stellar disks originate from the retrograde acquisition of gas, which settles on misaligned orbits with respect to the preexisting stellar disk and subsequently forms stars (Pizzella et al. 2004; Roškar et al. 2010; van de Voort et al. 2015; Starkenburg et al. 2019; Khoperskov et al. 2021, Bao et al. 2022; Bao et al. 2024; Gasymov et al. 2025). The CRDs are therefore direct evidence of past accretion events and constitute valuable tracers of galactic assembly. Several mechanisms have been proposed to explain the origin of the retrograde gas, including accretion from the cosmic web (Algorry et al. 2014), gas-rich mergers with dwarf galaxies (Thakar and Ryden 1996; Puerari and Pfenniger 2001; Algorry et al. 2014), tidal gas exchange between neighboring galaxies (Khim et al. 2021; Sil’chenko et al. 2023), and major mergers that preserve the stellar disk (Puerari and Pfenniger 2001; Crocker et al. 2009). Among these formation mechanisms, major mergers appear to be the least common in observed systems.
@@
Observationally, CRDs appear to be rare, with only 1% of late-type galaxies presenting them (Bao et al. 2022, Bevacqua et al. 2022, Gasymov et al. 2025). At the same time, CRDs are also challenging to observe due to their small mass contribution (Rubino et al. 2021). Nevertheless, modern integral field spectroscopy instruments can be used to observe these structures, which are typically identified from the presence of bimodal features in the velocity dispersion maps (Bao et al. 2022; Bao et al. 2024; Gasymov et al. 2025). Integral field spectroscopy instruments are also essential to disentangling the two stellar disk components and studying the kinematics and stellar population properties of CRDs in detail, which allows their formation scenarios to be constrained. Complementary techniques, such as spectral energy distribution fitting, are also used to extract physical and chemical properties of CRDs. Notable examples of CRDs have been identified in early-type disk galaxies. Examples are IC 719 (Katkov et al. 2013; Pizzella et al. 2018), NGC 4550 (Rubin et al. 1992; Rix et al. 1992; Coccato et al. 2013; Johnston et al. 2013), NGC 3593 (Bertola et al. 1996; Corsini et al. 1998; García-Burillo et al. 2000; Coccato et al. 2013), NGC 4191 (Coccato et al. 2015), NGC 448 (Katkov et al. 2016), NGC 1366 (Morelli et al. 2017), and NGC 5102 (Mitzkus et al. 2017). However, CRDs have also been observed in late-type disk galaxies, as reported in NGC 7217 (Merrifield and Kuijken, 1994), NGC 4138 (Jore et al. 1996; Pizzella et al. 2014), and NGC 5719 (Vergani et al. 2007; Coccato et al. 2011). These CRDs show a broad range of properties. They are generally observed to have younger stellar populations that corotate with the gas disk (Pizzella et al. 2004; Pizzella et al. 2014; Bao et al. 2022; Gasymov et al. 2025), steeper metallicity gradients, and broader metallicity distributions than the main stellar disk (Coccato et al. 2013; Bevacqua et al. 2022; Gasymov et al. 2025). As previously mentioned, the presence of CRDs is typically linked to past accretion events that induce star formation within the galaxy. Previous works have begun to explore the formation and evolution of CRD components using simulations (Lu et al. 2021; Santucci et al. 2024; Peirani et al. 2025). For example, using the IllustrisTNG100, Khoperskov et al. (2021) analyzed the properties and formation mechanisms of significant CRDs in early-type disk galaxies, finding that stellar counterrotation is mainly a result of an external gas infall. However, it is still unclear what the dominant CRD formation mechanisms are and under what conditions CRDs arise in late-type galaxies, topics that remain underexplored compared to early-type galaxy counterparts, where CRDs are more common (Kuijken et al. 1996; Bao et al. 2022; Bevacqua et al. 2022; Gasymov et al. 2025)
@@
In this work we focus on the search, characterization, and analysis of CRDs in Milky Way (MW)-type galaxies ($M_{\rm tot}\approx 10^{12}M_{\odot}$) in TNG50 from the IllustrisTNG suite of cosmological hydrodynamical simulations. We aim to determine how common CRDs are in MW-type galaxies as well as the dominant CRD formation mechanisms. In particular, we aim to confirm past formation channels proposed by observations of CRDs. The paper is organized as follows. In Section 2 we give an overview of the TNG50 simulation that we used, define our sample of galaxies, and provide our definition of counterrotation. In Section 3, an initial characterization of counterrotating components in the galaxies is shown, and afterward, the selection of significant CRDs is defined. In Section 4, the structural properties and assembly of the significant CRDs is shown, ending with a classification of the CRDs. We present analysis of the possible formation channels for the CRDs in each classification in Section 5. We discuss our results in Section 6. Finally, we provide a summary of our results and conclude in Section 7.
@@ 2 Methodology
In this section we discuss the simulations employed in this paper. We also describe the galaxy sample selection criteria and the definition of counterrotating stellar disks.
@@ 2.1 The IllustrisTNG: TNG50
Table 1: Parameters of the sample of galaxies with a significant CRD component and the six boundary cases. [^1] Column descriptions from the left to the right are: (1) internal name, (2) z=0 SubfindID from the TNG50 simulations, (3) total gravitationally bounded mass as defined by the Subfind algorithm, (4) stellar mass of the CRD, (5) total stellar disk mass, (6) half-mass radius of the stellar disk, (7) optical radius, (8) half-mass radius of the CRD, (9) radius enclosing 95% of the CRD, (10) stellar concentration of the CRD as defined in Sec. 4, (11) fraction of ex situ stellar particles in the CRD.
@@
The IllustrisTNG project (Nelson et al., 2015) is a set of gravitational magnetohydrodynamics cosmological simulations ran with the moving-mesh code AREPO (Springel, 2010). The project is made up of three simulation volumes: TNG50 (Nelson et al. 2019, Pillepich et al. 2019), TNG100, and TNG300 (Pillepich et al. 2018, Springel et al. 2018, Nelson et al. 2018, Naiman et al. 2018, Marinacci et al. 2018). Each of these runs solves the coupled evolution of dark matter, gas, stars, and supermassive black holes from $z=127$ to the present day, $z=0$. All of the simulations of the IllustrisTNG project are governed by a Lambda cold dark matter model with a cosmology consistent with the results of Planck Collaboration et al. (2016). The IllustrisTNG set of simulations has produced results consistent with observations in regimes beyond the ones used to calibrate the model (Nelson et al., 2015). For this work, we used the TNG50 run, and more specifically, we analyzed the baryonic simulation, TNG50-1 (TNG50, hereafter), which is the highest resolution simulation of the IllustrisTNG project. Its physical simulation box size is roughly 50 Mpc per side in length, and the stellar and dark matter particle mass resolutions are $8.5\times 10^{4}$ M and $4.5\times 10^{5}$ M, respectively. The softening lengths at $z=0$, are $\epsilon_{{\rm DM},*}=288$ pc and $\epsilon_{gas,min}=74$ pc. This simulation allows the components of galaxies to be analyzed in more detail than the other available simulations.
The TNG50 data provide the positions, velocities, mass, metallicity, and metal abundances of the stellar particles as part of the complete snapshot data. Each stellar particle represents a single stellar population with a given age and metallicity that is hereditary from the progenitor gas cell. Information about the galaxy and time where the stellar particle is born is provided in Rodriguez-Gomez et al. (2015) and is described in more detail in Section 2.3. The TNG50 simulation also gives information about the galaxies extracted from their member particles as a whole, classified according to a friends-of-friends (FoF) algorithm, and subsequently Subfind, including their total mass, total stellar mass, and virial radius. Information about the disk mass fraction according to the circularity of the particles is provided as an additional catalog by Genel et al. (2015).
@@ 2.2 Sample selection
In this study, we focus on a sample of MW-type galaxies; in a follow up study, we will expand this sample to include a wider range of late-type galaxies. In particular, we selected galaxies that fulfill the following criteria:
@@
  • The galaxy must be identified as the central galaxy of its FoF halo, defined as the most massive subhalo within that FoF group.
@@
In particular, the D/T ratios were obtained from the publicly available database of IllustrisTNG as part of the additional catalogs222www.tng-project.org/data/docs/specifications/#sec5c and were computed by considering as disk stellar particles the ones with a circularity parameter (see Sec 2.3) larger than 0.7.
@@
After applying the described selection criteria, we obtained a total of 260 late-type disk galaxies. As discussed, the selected galaxies have $\geq 5\times 10^{5}$ stellar particles within them, thus ensuring enough particles are present to reasonably resolve the structure of the stellar disk and search for potential structures with small mass contributions, such as counterrotating stellar components. The light blue histograms in Figure 1 show the distribution of stellar mass within the galactic disk (left panel, as defined in Sec. 2.3) and of the total mass (right panel) of the full sample of the full sample of 260 initially selected galaxies.
Figure 1: Distributions of mass of the full sample of galaxies and the galaxies hosting a significant CRD. Panel (a): Distribution of stellar mass confined within the disk region, independent of the circularity parameter. Panel (b): Distribution of the total mass according to the Subfind algorithm (i.e., considering all particles bound to the galaxy). In both panels, the distribution was normalized according to the full sample of 260 galaxies, represented by the light blue shaded distribution. The orange shaded region represents the distribution of the 26 galaxies hosting a significant CRD component.
@@
Figure 2: Distribution of the coplanar counterrotating disk component mass fraction for the full sample of 260 late-type disk galaxies. The vertical dashed orange line indicates a 1% mass fraction from the CRD. Only 10% of the galaxies (when including the boundary cases) have a CRD mass fraction greater than 1%.
Figure 3: Distribution of the circularity parameter for eight galaxies from the selected MW-type sample. The curves have been normalized so that the peak value is equivalent to one. The four top panels are galaxies that do not possess a significant CRD, while the bottom panels come from the subsample of 26 galaxies with a significant CRD. The shaded regions represent star particles identified as the CRD component (red) and the main stellar disk (light blue).
2.3 Coplanar counterrotating disk component definition
@@
The goal of this work is to identify, quantify, and analyze CRDs in fully cosmological simulations of galaxy formation. Different methods for isolating particles associated with different galaxy components, such as counterrotating disks, have been proposed in previous works. For example, Tissera et al. (2012) applied a criteria based on circularity and binding energy to subdivide a late-type galaxy based on its different components. In this study, we follow an approach more closely related to the definitions used in Gómez et al. (2017) and Monachesi et al. (2019). We first rotated the galaxy so that the plane of the disk is coplanar with the xy-plane, i.e., the angular momentum of the galactic disk is parallel to the z-axis direction. Afterward, we identified particles that are counter-rotating in the galaxy using the circularity parameter, thus performing a kinematic decomposition. Following Abadi et al. (2003), this parameter is defined as
@@
where $J_{z}$ is the angular momentum in the z-axis and the term $\mathrm{max}(J_{z}(E))$” indicates the maximum angular momentum in the z-direction allowed for that orbital energy, i.e., that of a circular orbit.
To define the extension of the stellar disks, we used the optical radius, $R_{\rm opt}$. As in Varela-Lavin et al. 2023 and Dolfi et al. 2023, $R_{\rm opt}$ is defined as the radius at which the surface brightness profile on the V-band of the galaxy, orientated in a face-on projection, reaches $\mu_{V}=25\ \mathrm{mag\ arcsec}^{-2}$. To estimate the optical radius, we measured the surface brightness of the galaxy in cylindrical shells of width 0.5kpc and height 10kpc above and below the disk plane. The values of the $R_{\rm opt}$ obtained for each galaxy are listed in Table 1.
Figure 4: Face-on and edge-on projections of the stellar mass surface density for a subsample of eight galaxies with the most significant counterrotating stellar disks. The galaxies are sorted according to their total mass, from the most to the least massive. We show all the stellar particles that belong to the host galaxy. The solid and dot-dashed yellow circles represent the radii, $R_{95}^{\rm CRD}$ and $R_{50}^{\rm CRD}$, of the CRD, respectively, while the dashed white circle represents the optical radius, $R_{\rm opt}$, of the galaxy.
Figure 5: Same as Fig. 4, but we show only the stellar particles belonging to the CRDs.
Having the circularity parameter computed for every stellar particle within $R_{\rm opt}$, we then defined the coplanar CRDs by selecting stellar particles that fulfill the following criteria:
@@
  • A distance from the galactic plane of less than 5kpc ($|z|<5\ \rm kpc$).
  • A galactocentric distance less than $R_{\rm opt}$.
These selection criteria allowed us to select star particles in nearly circular orbits rotating in the opposite direction of the overall disk. This definition is stricter than what has been used in past works, such as Khoperskov et al. (2021), and is closer to that of Jagvaral et al. (2022) and Lu et al. (2021) (more details in Appendix C).
In this work we are also interested on characterizing the formation mechanisms of CRDs. Due to this, it is important to identify the origins of the different stellar particles associated to these components. In particular, we subdivide the counterrotating stellar particles as in situ or ex situ based on where the particle were born. As discussed in Rodriguez-Gomez et al. 2016, a particle is considered in situ if the galaxy in which it formed is part of the main branch of the merger tree of the galaxy where the particle is at the present-day. In case the particle was born outside this main branch then the particle is tagged as ex situ. Notice that particles that are formed within the potential well of the main progenitor from recently stripped gas from satellite galaxies are labeled as in situ.
3 Identification of coplanar counterrotating disk components